If the tangents at two points of a parabola are at right angles, then show that they intersect at a point on the directrix.
1
Expert's answer
2020-06-24T18:09:20-0400

We remind that the parabola is the locus of points that are equidistant from both focus and the directrix . We assume that the focus has coordinates "(x_0,y_0)" and "y=y_1" is an equation of directrix. The following equality has to be satisfied:

"|y-y_1|=\sqrt{(x-x_0)^2+(y-y_0)^2}"

From the latter, by direct calculations, we get:

"(y-y_1)^2=(x-x_0)^2+(y-y_0)^2\,\,\,\Longrightarrow\,\,\,\,y=\frac{(x-x_0)^2}{2(y_0-y_1)}" .

We remind that the equation of the tangent line to the curve "y=f(x)" (see e.g., https://en.wikipedia.org/wiki/Tangent) at point "\tilde{x}" has the form:

"y=f(\tilde{x})+f'(\tilde{x})(x-\tilde{x})"

From this we receive that equation of the tangent line to parabola has the form:

"y=\frac{(\tilde{x}-x_0)^2}{2(y_0-y_1)}+\frac{\tilde{x}}{y_0-y_1}(x-\tilde{x})"

Suppose that we have two tangent lines at points "\tilde{x}_1" and "\tilde{x}_2". The tangent of the angle between those two lines is (see https://en.wikipedia.org/wiki/List_of_trigonometric_identities):

"tan\,\alpha=\frac{\frac{\tilde{x}_1-\tilde{x}_2}{y_0-y_1}}{1+\frac{\tilde{x}_2}{y_0-y_1}\cdot\frac{\tilde{x}_1}{y_0-y_1}}=\frac{\tilde{x}_1-\tilde{x}_2}{y_0-y_1+{\tilde{x}_2}\tilde{x}_1}"

The tangent tends to infinity for the angles close to 900

This means, that "y_0-y_1+{\tilde{x}_2}\tilde{x}_1=0". From the latter we found that "\tilde{x}_2=\frac{y_1-y_0}{\tilde{x}_1}". In order to find the point, where two lines intersect, we have the equality:

"\frac{(\tilde{x}_1-x_0)^2}{2(y_0-y_1)}+\frac{\tilde{x}_1}{y_0-y_1}(x-\tilde{x}_1)=\frac{(\tilde{x}_2-x_0)^2}{2(y_0-y_1)}+\frac{\tilde{x}_2}{y_0-y_1}(x-\tilde{x}_2)"

After simplifications we receive:

"x=x_0+\frac{\tilde{x}_1+\tilde{x}_2}{2}"

We substitute the latter in equation for the tangent line and receive the y-coordinate of the point, where tangent lines intersect. We point out, that in general it is not equal to "y_1"